4) Answer the following a) How many moles of H were added to the buffers by adding 0.50 mL of 1.0 M HCl? a b) How many moles of conjugate base (A-) are present in 40 mL of the 0.10 M Buffer at pH 7.5? c) How many mol es of conjugate base (A-) are present in 40 mL of the 0.10 M Buffer at pH 6.0? d) Use the Henderson-Hasselbalch equation to predict how much the pH should change (i.e. ApH) when 0.50 mL of 1.0 M HCI are added to 40 mL of the 0.10 M buffer at pH 6.0 vs. pH 7.5. 7.5 vs. at pH 6.0? of moles of conjugate base present in 40 mL of the 0.010 M buffers at each pH? e) How many moles of conjugate base (A-) are present in 40 mL of the 0.010 M Buffer at pH f) How do the number of moles of H that are added in this experiment compare to the number
Answer #1

A) moles of H+ = moles of HCl = Molarity × volume in ml/1000

= 1× 0.5 ml/ 1000 = 5 × 10-4 moles

B) Sorry I can't further do because neither you have given pKa or nor which buffer you are working.

pH = pKa + Log10( [A-]/[HA])

From above equation pKa is required.